OpenStudy (anonymous):

tree that is 100 feet tall casts a shadow that is 150 feet long. Determine the angle at which the rays of the sun hit the ground, to the nearest degree.

OpenStudy (anonymous):



OpenStudy (anonymous):

tan of anglle = 100/150 = 2/3

OpenStudy (anonymous):

is it 42 degrees?